OpenStudy (anonymous):

write the equation of the line that passes through (2,1) and (5,-5)

OpenStudy (aroub):

first find the slope : -5-1/ 5-2 = -6/3 = -2 y=mx+b , m=-2 , (2,1) 1=-2(2)+b 1=-4+b 4+1=b 5=b so the equation is y=-2x+5